Find LCD for list of rational expressions, 13/8x , 7/4x+8
Answer #1

$$ \frac{13}{8 x}, \frac{7}{4 x+8} $$

for \(L C D\)

Note down the denominator of each fraction.

$$ 8 x, 4 x+8 $$

find the LCM of the denominators

$$ \begin{aligned} &8 x, 4(x+2) \\ &8 x=2 \times 2 \times 2 \times x \\ &4(x+2)=2 \times 2 \times(x+2) \end{aligned} $$

now LCM of denominators

$$ \begin{aligned} &=2 \times 2 \times 2 \times x \times(x+2) \\ &=8 x(x+2) \end{aligned} $$

so

$$ L C D=8 x(x+2) $$